# Projects/QSM24/SFVisco/ViscoSF.tex

### v4

@@ II.2 Adiabatic Transition
$\approx$ $\frac{\mbox{sign}(\tau F_{{\bf a},\alpha})\tilde{\Delta}_{p}{a}_{\alpha}}{m}\wp^{+}_{\alpha}({\bf a}),$
again to leading order. Since this must equal the negative of the change in potential energy in the $\alpha$ direction, $({\tau}/{m}){F}_{{\bf a},\alpha}a_{\alpha}$, we obtain
$\wp^{+}_{\alpha}({\bf a})=\frac{|\tau{F}_{{\bf a},\alpha}|}{\tilde{\Delta}_{p}}.$ (2.19)
This must be positive and less than unity, which can be ensured by choosing $\tilde{\Delta}_{p}=n_{\alpha}\Delta_{p}$, where $n_{\alpha}$ is a positive integer. Choose $n_{\alpha}$ to be as small as possible, but no smaller. Practical experience with the finite-sized simulations reported below has shown unity to be sufficient, $\tilde{\Delta}_{p}=\Delta_{p}$. The smaller the time step the smaller $\wp^{+}_{\alpha}({\bf a})$ is and the more likely the boson is to remain in its current momentum state.
Requiring the cancelation of kinetic and potential energies on average ensures energy conservation and the constancy of the equilibrium probability distribution on the stochastic adiabatic trajectory. It means that on average the change in momentum of the bosons in the state is equal to the classical change in momentum due to the force acting on them. Hence we expect that that these stochastic equations of motion will go over to the classical deterministic equations of motion in the thermodynamic limit in the non-condensed regime.
II.3 Dissipative Transition
@@
The dissipative transitions act like a thermostat and provide a more direct mechanism for the change in occupancy of the momentum states and for the equilibration of the occupancy distribution. For this we first randomly choose a boson from the $N$ bosons in the subsystem, say $j$. We then choose with probability in inverse proportion to the occupancy of its state whether or not this boson will attempt a transition. (Attempt a transition only if a random number uniformly distributed on the unit interval is less than $1/N_{{\bf p}_{j}}$.) This step is necessary so that all occupied states are treated equally irrespective of their occupancy. Finally, if a transition for $j$ is to be attempted we use the conditional transition probability that follows for the 27 near neighbor states ${\bf a}^{\prime}$ (including the original state ${\bf a}$). Typically, this sequence of steps is repeated in a block of $N$. The block is performed typically once every 10 time steps, although less frequent attempts would probably suffice.
The irreversible form of Bayes’ theorem is required,
$\frac{\wp(N_{{\bf a}^{\prime}}+1,N_{{\bf a}}-1|N_{{\bf a}^{\prime}},N_{{\bf a}})}{\wp(N_{{\bf a}^{\prime}},N_{{\bf a}}|N_{{\bf a}^{\prime}}+1,N_{{\bf a}}-1)}$ $=$ $\frac{\wp(N_{{\bf a}^{\prime}}+1)\wp(N_{{\bf a}}-1)}{\wp(N_{{\bf a}^{\prime}})\wp(N_{{\bf a}})}$ (2.20)
$=$ $e^{-\beta(a^{\prime 2}-a^{2})/2m}.$
There is no time step involved in this. Now ${\bf a}^{\prime}={\bf a}+{\bf s}\Delta_{p}$ and $a^{\prime 2}-a^{2}=2\Delta_{p}{\bf s}\cdot{\bf a}+\Delta_{p}^{2}s^{2}$, where $s_{\alpha}\in\{-1,0,1\}$. In view of classical dissipative dynamics based on the second entropy (Attard 2012), the conditional transition probability for ${\bf a}\stackrel{{\scriptstyle j}}{{\to}}{\bf a}^{\prime}$ has the form
$\wp({\bf a}^{\prime}|{\bf a})=\wp_{0}+({\bf a}^{\prime}-{\bf a})\cdot{\bf R}_{p}({\bf a})/\Delta_{p}^{2}.$ (2.21)
Normalization gives $\wp_{0}=1/27$ and
${\bf R}_{p}({\bf a})=\frac{-\Delta_{p}^{2}}{54mk_{\mathrm{B}}T}{\bf a}.$ (2.22)
It can be confirmed that with this Bayes’ theorem is satisfied to quadratic order. This is just the quantized version of the classical drag force.
The heart of the dissipative part of the algorithm is to calculate $\wp({\bf a}^{\prime}|{\bf a})$ for each of the 27 nearest neighbors (including ${\bf a}$), and place these in some one-dimensional order. This divides the unit interval into subintervals of width equal to the respective transition probability. Next choose a random number uniformly distributed on the unit interval. The boson $j$ is moved to the momentum state corresponding to the subinterval into which the random number falls. Obviously if this subinterval corresponds to the original state ${\bf a}$, then no transition is made. The procedure is repeated for a new randomly chosen boson. There is nothing to prevent a boson being chosen more than once in the current block, in which case it could make several transitions, while other bosons have not been chosen at all.
This random choice of $N$ bosons for transition is not essential. It was used for the quantum fluid for the results below. In what is called the classical fluid, all $N$ bosons were cycled through in sequential order for an attempted transition. This difference between the treatment of the two fluids is not expected to be significant. For the quantum fluid a transition was attempted with probability in inverse proportion to the occupancy of its state. For the classical fluid, a transition was always attempted irrespective of the occupancy of its momentum state. This difference is significant.
This quantum dissipative algorithm was tested for ideal bosons. It was found that for $N=1,000$, $T^{*}=0.60$, $\rho^{*}=0.80$, it gave the ground state occupancy as $\langle N_{000}\rangle=22.8(9)$, which can be compared to the exact result 38.9. The results for the excited states were better; for example the first excited state had $\langle N_{001}\rangle=10.9(2)$ compared to the exact 12.5. For $N=10,000$ and the same density and temperature, the algorithm gave $\langle N_{000}\rangle=35.1(38)$, with the exact result in this case being 40.9. The first excited state was $\langle N_{001}\rangle=26.2(24)$, compared with the exact 27.9. This dependence on system size for the ground state occupancy is acceptable for the present purposes.
II.3.1 Alternative Expansion
As an alternative, expand the right hand side as
$\frac{\wp(N_{{\bf a}^{\prime}}+1)\wp(N_{{\bf a}}-1)}{\wp(N_{{\bf a}^{\prime}})\wp(N_{{\bf a}})}$
$=$ $e^{-\beta(a^{\prime 2}-a^{2})/2m}$
@@ II.3.1 Alternative Expansion
$=$ $1-\frac{\beta}{2m}(a^{\prime 2}-a^{2})+\frac{\beta^{2}}{8m^{2}}(a^{\prime 2}-a^{2})^{2}+{\cal O}(\Delta_{p}^{3}).$
For the conditional transition probability use the expansion
$\wp(N_{{\bf a}^{\prime}}+1,N_{{\bf a}}-1|N_{{\bf a}^{\prime}},N_{{\bf a}})$
$=$ $A({\bf a})+B({\bf a})(a^{\prime 2}-a^{2})+C({\bf a})(a^{\prime 2}-a^{2})^{2}+{\cal O}(\Delta_{p}^{3}).$
It is straightforward to show that
$A({\bf a})=\frac{1}{27}+{\cal O}(\Delta_{p}^{2}),\mbox{ and }B({\bf a})=\frac{-\beta}{108m}+{\cal O}(\Delta_{p}^{2}).$ (2.25)
The coefficient $C({\bf a})$ can be set to zero and the expansion of the right hand side will remain satisfied at ${\cal O}(\Delta_{p}^{2})$.
Results for ideal bosons ($N=1,000$, $T=0.60$, $\rho=0.80$) gave the ground state occupancy as $\langle N_{000}\rangle=38.0(8)$, compared to the exact 38.9, and the first excited state occupancy as $\langle N_{001}\rangle=13.3(1)$, compared to the exact 12.5. For $N=10,000$ the algorithm gave $\langle N_{000}\rangle=41.9(40)$, compared to the exact 40.9, and $\langle N_{001}\rangle=28.4(9)$, compared to the exact 27.9.
Although these results for ideal bosons are better than those obtained with the second entropy algorithm, preliminary results suggested that the latter was better for interacting bosons, as judged by the kinetic energy in the classical case. Also the second entropy approach has fundamental justification in the classical limit (Attard 2012a). For these reasons it is the second entropy algorithm that is used to obtain the results given below.
II.4 Viscosity Time Correlation Function
@@
$P_{\alpha\gamma}=\sum_{j=1}^{N}q_{j,\alpha}p_{j\gamma}.$ (2.27)
@@
This can be shown to be symmetric, $\dot{P}^{0}_{\gamma\alpha}=\dot{P}^{0}_{\alpha\gamma}$.
In the occupancy picture we cannot distinguish individual bosons, and so we should replace here the classical force on individual bosons by the shared non-local force on the bosons in each momentum state. Arguably this is equivalent to using the adiabatic stochastic transition probability since on average the two are the same. Therefore, for the momentum state ${\bf a}$, define the shared non-local force per boson and the center of mass,
${\bf F}_{\bf a}\equiv\frac{1}{N_{\bf a}}\sum_{j\in{\bf a}}{\bf f}_{j},\mbox{ and }{\bf Q}_{\bf a}\equiv\frac{1}{N_{\bf a}}\sum_{j\in{\bf a}}{\bf q}_{j}.$ (2.29)
Also define the total force on the momentum state ${\bf a}$ due to the state ${\bf b}$ to be
${\bf F}_{{\bf a}{\bf b}}=\sum_{j\in{\bf a}}\sum_{k\in{\bf b}}{\bf f}_{jk},$ (2.30)
where ${\bf f}_{jk}$ is the classical force on boson $j$ due to boson $k$. Note that ${\bf F}_{\bf a}=N_{\bf a}^{-1}\sum_{\bf b}{\bf F}_{{\bf a}{\bf b}}$. According to Newton, ${\bf F}_{{\bf a}{\bf b}}=-{\bf F}_{{\bf b}{\bf a}}$ and ${\bf F}_{{\bf a}{\bf a}}={\bf 0}$. Using these the adiabatic rate of change of the first momentum moment dyadic in the condensed regime is
$\dot{\underline{\underline{P}}}^{0}$ $=$ $\frac{1}{m}\sum_{j=1}^{N}{\bf p}_{j}{\bf p}_{j}+\sum_{{\bf a}}N_{\bf a}{\bf Q}_{{\bf a}}{\bf F}_{{\bf a}}$ (2.31)
$=$ $\frac{1}{m}\sum_{j=1}^{N}{\bf p}_{j}{\bf p}_{j}+\sum_{{\bf a}}{\bf Q}_{{\bf a}}\sum_{\bf b}{\bf F}_{{\bf a}{\bf b}}$
$=$ $\frac{1}{m}\sum_{j=1}^{N}{\bf p}_{j}{\bf p}_{j}+\frac{1}{2}\sum_{{\bf a},{\bf b}}{\bf Q}_{{\bf a}{\bf b}}{\bf F}_{{\bf a}{\bf b}},$
where ${\bf Q}_{{\bf a}{\bf b}}={\bf Q}_{{\bf a}}-{\bf Q}_{{\bf b}}$. For the shared non-local force, ${\bf f}_{j}\Rightarrow{\bf F}_{{\bf p}_{j}}$, which means that the transformation to the center of mass is exact because all the bosons in the state experience the same force, $\sum_{j\in{\bf a}}{\bf q}_{j}{\bf f}_{j}=\sum_{j\in{\bf a}}{\bf q}_{j}{\bf F}_{{\bf a}}=N_{\bf a}{\bf Q}_{\bf a}{\bf F}_{{\bf a}}$. In so far as the momentum state is non-local, there are no spatial correlations between the bosons in the state, and on average ${\bf Q}_{\bf a}\approx{\bf 0}$ (although the fluctuations in this are macroscopic). Similarly the shared non-local force likely averages close to zero for highly occupied states (again with large fluctuations). These two properties doubtless contribute to the molecular origin of the vanishing of the viscosity for condensed bosons.
Numerical results indicate that this dyadic is symmetric on average, in the same sense that all off-diagonal components of the dyadic are equal on average. The standard deviation of the six off-diagonal components of the viscosity time function gives an estimate of the statistical error.
@@ III Results
In what follows what is defined as the classical fluid is treated with the same algorithm as the quantum fluid but as if the occupancy of the momentum state was unity. That is, the force on each boson was taken to be the actual local force rather than the shared non-local force, and the dissipative transition was always attempted for the chosen boson irrespective of the occupancy of its momentum state. The viscosity was calculated with the actual local force rather than the shared non-local force. However, the momenta were still quantized and the adiabatic transitions were still stochastic, just as in the quantum case.
We used the Lennard-Jones pair potential,
@@
We set this to zero for $r>3.5\sigma$. The molecular diameter for helium is $\sigma_{\mathrm{He}}=0.2556$ nm and the well-depth is $\varepsilon_{\mathrm{He}}/k_{\mathrm{B}}=10.22$ J (van Sciver 2012). The mass is $m_{\mathrm{He}}=4.003\times 1.66054\times 10^{-27}$ Kg.
The results below are presented in dimensionless form: the temperature is $T^{*}=k_{\mathrm{B}}T/\epsilon_{\mathrm{He}}$, the number density is $\rho^{*}=\rho\sigma_{\mathrm{He}}^{3}$, the time step is $\tau^{*}=\tau/t_{\mathrm{He}}$, the unit of time is $t_{\mathrm{He}}=\sqrt{m_{\mathrm{He}}\sigma_{\mathrm{He}}^{2}/\epsilon_{\mathrm{He}}}$, and the shear viscosity is $\eta^{*}=\eta\sigma_{\mathrm{He}}^{3}/\epsilon_{\mathrm{He}}t_{\mathrm{He}}$.
A canonical system with $N=1,000$ Lennard-Jones atoms was simulated with the above quantum stochastic molecular dynamics algorithm. The time step was $\tau^{*}=2\times 10^{-5}$. A block of $N$ dissipative transitions was attempted once every 10 time steps. For the classical liquid this combination consistently gave a kinetic energy 2–3% higher than the exact value. The present parameters were judged acceptable even though a smaller time step or more frequent dissipative transitions would have given more accurate results, albeit at the cost of longer simulation times or greater interference in the adiabatic evolution of the momentum moment. Averages were collected once every 75 time steps. The number of time steps in a run was $10\times 4,000\times 75$. A run took about 24 hours on a desk-top personal computer. Four runs were made at each thermodynamic point. The quoted statistical error is the larger of the error estimated from the fluctuations in 10 blocks of each run (averaged across the runs), or the standard deviation across the four runs. There was usually little difference in the two estimates (but see below). The shear viscosity time function was estimated for $t^{*}\leq 6$, which meant that the time correlation function was covered ten times, and that $4\times 10\times 6$ values were averaged at each time point for the quantum case, and $4\times 10\times 3$ values in the classical case. The components of the quantized momentum were restricted to $p_{j\alpha}^{2}/2mk_{\mathrm{B}}T\leq 13$.
@@
Figure 2: Ground state occupancy in the saturated Lennard-Jones liquid. The circles are the quantum liquid, the triangles are the classical liquid, and the crosses are the exact result for ideal bosons. The error bars are less than the symbol size. The lines are an eye guide.
Figure 2 shows the ground state occupancy on the saturation curve for the Lennard-Jones liquid. It can be seen that the present quantum molecular dynamics algorithm for the quantum liquid gives a larger occupancy than for the classical liquid, but smaller than that for ideal bosons. It has been argued that the ideal boson result should approximately apply to interacting bosons on the far side of the $\lambda$-transition (Attard 2025 §5.3). The smaller value evident in the figure could be due to finite size effects, or it could be due to forces in the interacting liquid pushing the ground state to neighboring momentum states, or it could be that the ideal boson model is a poor approximation for interacting bosons.
We defined a quantity maxocc to be the occupancy of the maximally occupied momentum state at each instant. For $T^{*}=0.65$, ${\tt maxocc}=22.3(2)$ for the quantum liquid, and 5.5773(4) for the classical liquid. (Here and throughout, the quantity in parentheses is the statistical error for the final digit or digits. It gives the 95% confidence level, as is also the case for the error bars in the figures.) The quantum result is close to the ideal boson result for the ground state occupancy, $N_{000}^{\mathrm{id}}=29.3$. However, at the higher temperature $T^{*}=0.90$, ${\tt maxocc}=9.71(3)$ in the quantum case, compared to $N_{000}^{\mathrm{id}}=2.69$. The ground state occupancy for the quantum liquid, $N_{000}^{\mathrm{qu}}=2.14(3)$ is not too bad at this temperature.
In any case the occupancy of the ground state is clearly greater for the quantum liquid than for the classical liquid. However, the occupancy of the ground state in the quantum liquid is a small fraction of the total number of bosons in the subsystem. For example, the fraction of the subsystem in the ground state is $1\%$ at $T^{*}=0.65$. Obviously this means that ground state condensation cannot account for the $\lambda$-transition or for superfluidity.
@@
Figure 3 gives the fraction of condensed bosons in the subsystem. Condensed bosons were defined as those in multiply occupied states, with a threshold set at 2 or 3. In the quantum liquid at the lowest temperature studied about 69% of the bosons are in states with two or more, and about 53% are in states with three or more. In contrast the fraction for the classical liquid is 47% and 20%, respectively. We can conclude that Bose-Einstein condensation is substantial, and that multiple momentum states are multiply occupied. That the majority of the bosons in the system can be considered to be condensed explains the macroscopic nature of the $\lambda$-transition and superfluidity.
It can be seen that at higher temperatures the condensation in the quantum liquid is approaching that in the classical liquid, as one would hope. However even at the highest temperature studied, $T^{*}=1$, there is still excess condensation in the quantum liquid, $f_{0}^{\mathrm{qu}}=40\%$ compared to $f_{0}^{\mathrm{cl}}=23\%$. That there remains condensation well-above the superfluid transition temperature is in part due to the neglect in the present calculations of position permutation loops, which suppress condensation (Attard 2025 §3.2). This has implications for the two-fluid model of superfluidity, which is discussed below.
The kinetic energy per boson in the classical liquid is $\beta{\cal K}/N=1.5406(6)$ at $T^{*}=0.90$ and 1.5366(3) at $T^{*}=0.60$. The equipartition theorem gives the exact classical value as 3/2. One can conclude that the present stochastic equations of motion that use the transition probability for quantized momentum are indeed going over to the classical equations of motion. The kinetic energy per boson in the quantum liquid is $\beta{\cal K}/N=1.2859(8)$ at $T^{*}=0.90$ and 1.001(1) at $T^{*}=0.60$. The decrease in kinetic energy with decreasing temperature is a manifestation of the increasing condensation in the quantum liquid that preferentially occurs in the low lying momentum states.
@@
Figure 4 shows the shear viscosity time function for the quantum and classical liquids at the relatively high temperature of $T^{*}=1.00$. The general features of the curves is that they rise from zero at $t^{*}=0$ to reach a maximum or a plateau, in this case at about $t^{*}=6$. This maximum value is taken to be ‘the’ shear viscosity. It can be seen that the viscosity of the classical liquid is greater than that of the quantum liquid, although they do agree to within the statistical error over most of the time domain shown. The fraction of condensed bosons in states with two or more bosons at this temperature is 40% for the quantum liquid and 23% for the classical liquid (Fig. 3). Hence although one expects the quantum viscosity to go over to the classical viscosity as the temperature is increased, there still remains substantial Bose-Einstein condensation at this temperature.
The diffusion in the system was relatively small. The root mean square change in separation of a randomly chosen pair of bosons over the course of a run was 0.91(32) for the quantum liquid and 0.78(31) for the classical liquid. Nevertheless the statistical errors in the average potential energy and pressure were about the same whether estimated within a run or between runs. (Each run was started from an independently equilibrated configuration.) I conclude from this that the subsystem is liquid-like rather than glassy.

### v5

@@ II.4 Distribution of Subset Size
$=$ $\frac{|\tau f_{j,\alpha}|}{N_{{\bf a}}\Delta_{p}}\left\{1-\frac{\beta\Delta_{p}}{2m}\mbox{sign}(\tau{f}_{j\alpha})a_{\alpha}+\frac{\beta\tau a_{\alpha}}{2m}{f}_{j\alpha}\right\}.$
Here ${\bf a}^{\prime}_{j\alpha}={\bf a}+\mbox{sign}(\tau f_{j\alpha})\Delta_{p}\widehat{\bf x}_{\alpha}$. Sequential transitions for each component and for each boson in the momentum state are attempted at each time step. There appears to be little difference in whether the occupancy is updated after each successful transition, or only at the end of the time step. There also appears little difference if the quadratic term in the change in kinetic energy is added. And it is also possible to use the exponential form of the term in braces.
The restriction to individual transitions, $g(n_{A})=\delta_{n_{A},1}$, means that the damping of the rate of change of momentum is polynomial rather than exponential, $\propto N_{\bf a}^{-1}$. Since the occupancy of low-lying momentum states can by ${\cal O}(10^{2})$ below the condensation transition, this is still a significant reduction, and it appears sufficient to be interpreted as the origin of superfluidity.
II.5 Dissipative Transition
@@
The dissipative transitions act like a thermostat and provide another mechanism for the change in occupancy of the momentum states and for the equilibration of the occupancy distribution. For this I randomly or sequentially choose a boson from the $N$ bosons in the subsystem, say $j$. I use the following conditional transition probability for the 27 near neighbor states ${\bf a}^{\prime}$ (including the original state ${\bf a}$). The block of $N$ individual attempted transitions is performed typically once every 10 time steps, although less frequent attempts would probably suffice.
The dissipative transition is irreversible, which means that the forward and backward unconditional transitions are equally likely. Hence for the transition to a neighboring momentum state ${\bf a}\stackrel{{\scriptstyle j}}{{\to}}{\bf a}^{\prime}$, the ratio of conditional transition probabilities is
$\frac{\wp_{j}({\bf a}^{\prime}|{\bf a})}{\wp_{j}({\bf a}|{\bf a}^{\prime})}$ $=$ $\frac{\wp_{j}({\bf a}^{\prime})}{\wp_{j}({\bf a})}$ (2.20)
$=$ $\frac{N_{{\bf a}^{\prime}}+1}{N_{\bf a}}\left[1-\frac{\beta(a^{\prime 2}-a^{2})}{2m}\right].$
Here the change in kinetic energy has been expanded to quadratic order. This is satisfied by
@@
$\wp_{j}({\bf a}^{\prime}|{\bf a})=\left\{\begin{array}[]{ll}\displaystyle\frac{\varepsilon}{N_{\bf a}}\left[1-\frac{\beta(a^{\prime 2}-a^{2})}{4m}\right],&{\bf a}^{\prime}\neq{\bf a}\\ \displaystyle 1-\frac{26\varepsilon}{N_{\bf a}}+\frac{54\beta\Delta_{p}^{2}}{4m}\frac{\varepsilon}{N_{\bf a}},&{\bf a}^{\prime}={\bf a}.\end{array}\right.$ (2.21)
For the following results, $\varepsilon=1/27$. The dissipative transitions were attempted one boson at a time, for all $N$ bosons in a cycle. A cycle of such attempts was made once every skipcon adiabatic time steps. These dissipative transitions roughly correspond to $r=1$, $g(n_{A})=\delta_{n_{A},1}$, and $\lambda=1/{\tt skipcon}$.
The present algorithm has proven adequate to ensure the equilibrium distribution, although it is not actually clear that a dissipative thermostat is required because unlike the classical adiabatic equations of motion, temperature already appears in the present adiabatic conditional transition probability.
II.5.1 Ideal Boson Results
For ideal bosons, this second order dissipative transition algorithm for $T=0.60$, $\rho=0.8872$, and $N=1,000$, gave for the occupancies of the first several low lying momentum states: $N_{000}=85.9(43)$, (exact is 87.0), $N_{001}=16.1(3)$, (exact is 14.5), $N_{011}=8.2(1)$, (exact is 7.73), $N_{111}=5.39(4)$, (exact is 5.17), $N_{002}=3.94(3)$, (exact is 3.83).
For $T=0.60$, $\rho=0.8872$, and $N=10,000$, the results were $N_{000}=138.2(284)$, (exact is 138.0), $N_{001}=52.0(31)$, (exact is 51.9), $N_{011}=33.3(8)$, (exact is 31.8), $N_{111}=23.5(5)$, (exact is 22.8), $N_{002}=18.3(6)$, (exact is 17.8). These improve significantly the results of the linear algorithm given in earlier versions of this paper.
II.6 Viscosity Time Correlation Function
@@
$P_{\alpha\gamma}=\sum_{j=1}^{N}q_{j\alpha}p_{j\gamma}.$ (2.23)
@@
The force ${\bf f}_{j}$ gives the classical rate of change of momentum of boson $j$. This can be cast in symmetric form using the gradient of the pair potential, in which form the minimum image convention can be applied.
In the condensed regime, the average rate of change of momentum for boson $j$ to leading order is
$\left\langle\dot{\bf p}_{j}^{0}\right\rangle=\frac{1}{N_{{\bf p}_{j}}}{\bf f}_{j}.$ (2.25)
This uses $g(n_{A})=\delta_{n_{A},1}$, Eq. (II.4). With this the adiabatic rate of change of the first momentum moment is
$\dot{\underline{\underline{P}}}^{0}$ $=$ $\frac{1}{m}\sum_{j=1}^{N}{\bf p}_{j}{\bf p}_{j}+\sum_{j=1}^{N}{\bf q}_{j}\frac{1}{N_{{\bf p}_{j}}}{\bf f}_{j}$ (2.26)
$=$ $\frac{1}{m}\sum_{j=1}^{N}{\bf p}_{j}{\bf p}_{j}+\frac{1}{2}\sum_{j,k}\tilde{\bf q}_{jk}{\bf f}_{j,k}.$
Here ${\bf f}_{j,k}$ is the force on boson $j$ due to boson $k$, so that the total force on boson $j$ is ${\bf f}_{j}=\sum_{k}{\bf f}_{j,k}$, and
$\tilde{\bf q}_{jk}\equiv\frac{1}{N_{{\bf p}_{j}}}{\bf q}_{j}-\frac{1}{N_{{\bf p}_{k}}}{\bf q}_{k}.$ (2.27)
Generally simulations are conducted with periodic boundary conditions, which give a homogeneous system. In the uncondensed regime, $N_{\bf a}=1$ and $\tilde{\bf q}_{jk}={\bf q}_{jk}$. The minimum image separation is ${\bf q}_{jk}={\bf q}_{jk}-L*{\tt ANINT}({\bf q}_{jk}/L)$, where ANINT gives the nearest integer. This guarantees that $|q_{jk,\alpha}|\leq L/2$ and it is known to work in the classical case (ie. it gives a viscosity that is independent of system size). In the condensed regime, $N_{{\bf p}_{j}}\geq 1$, and so the minimum image convention can still be applied to $\tilde{\bf q}_{jk}$ with the same guarantee.
In the simulations, the adiabatic rate of change of the first momentum moment was calculated once every 75 time steps. The time correlation function was constructed and integrated to give the viscosity time function. This was averaged over the six components (three in the classical case) and the maximum value taken to be ‘the’ shear viscosity.
@@ III Results
In what follows what is defined as the classical fluid is treated with the same algorithm as the quantum fluid but as if the momentum states were solely occupied. Hence individual transitions were attempted without the factor of $N_{\bf a}^{-1}$ in the conditional probability or in the rate of change of the first momentum moment in the shear viscosity. However, the momenta were still quantized and the adiabatic transitions were still stochastic, just as in the quantum case.
The Lennard-Jones pair potential was used,
@@
This was set to zero for $r>3.5\sigma$. The molecular diameter for helium is $\sigma_{\mathrm{He}}=0.2556$ nm and the well-depth is $\varepsilon_{\mathrm{He}}/k_{\mathrm{B}}=10.22$ J (van Sciver 2012). The mass is $m_{\mathrm{He}}=4.003\times 1.66054\times 10^{-27}$ kg.
The results below are presented in dimensionless form: the temperature is $T^{*}=k_{\mathrm{B}}T/\varepsilon_{\mathrm{He}}$, the number density is $\rho^{*}=\rho\sigma_{\mathrm{He}}^{3}$, the time step is $\tau^{*}=\tau/t_{\mathrm{He}}$, the unit of time is $t_{\mathrm{He}}=\sqrt{m_{\mathrm{He}}\sigma_{\mathrm{He}}^{2}/\varepsilon_{\mathrm{He}}}$, and the shear viscosity is $\eta^{*}=\eta\sigma_{\mathrm{He}}^{3}/\varepsilon_{\mathrm{He}}t_{\mathrm{He}}$.
A canonical system, most commonly with $N=1,000$ Lennard-Jones atoms, was simulated with the above quantum molecular dynamics algorithm. The time step was $\tau^{*}=2\times 10^{-5}$. A block of $N$ dissipative transitions was attempted once every 10 time steps. For the classical liquid this combination consistently gave a kinetic energy 5-7% lower than the exact value. The present parameters were judged acceptable even though a smaller time step or a larger system would have given more accurate results albeit at the cost of longer simulation times. More frequent dissipative transitions would also likely improve the classical kinetic temperature but at the cost of greater interference in the adiabatic evolution of the momentum moment. Averages were collected once every 75 time steps. The number of time steps in a run was $10\times 4,000\times 75$. A run took 24–30 hours on a desk-top personal computer. Typically 4–8 runs were made at each thermodynamic point, although twelve were made in the classical cases at the lowest temperatures. The quoted statistical error is the larger of the error estimated from the fluctuations in 10 blocks of each run (averaged across the runs), or the standard deviation across the runs. There was usually little difference in the two estimates (but see below). The shear viscosity time function was estimated for $t^{*}\leq 6$, which meant that the time correlation function was covered ten times, and that in a single run $10\times 6$ values were averaged at each time point for the quantum case, and $10\times 3$ values in the classical case. In the quantum case the viscosity time function had more or less reached it maximum by $t=6$. In the classical case the viscosity time function was extrapolated beyond $t=6$ to its maximum value using a quadratic best fit. The quantum viscosity time function was rather flat and no extrapolation was made. The components of the quantized momentum were restricted to $p_{j\alpha}^{2}/2mk_{\mathrm{B}}T\lesssim 20$.
@@
Figure 1: Ground state occupancy in the saturated Lennard-Jones liquid. The circles are the quantum liquid, the triangles are the classical liquid, and the crosses are the exact result for ideal bosons. The error bars are less than the symbol size. The lines are an eye guide. Note that $T[K]=10.22T^{*}$.
Figure 1 shows the ground state occupancy on the saturation curve for the Lennard-Jones liquid. It can be seen that the simulated occupancy in the quantum liquid is in good agreement with the analytic result calculated for non-interacting bosons. Comparable if not better agreement holds for the first several excited momentum states. The ideal boson result should apply to interacting bosons on the far side of the $\lambda$-transition (Attard 2025 §5.3). The slightly larger than ideal value for the quantum liquid ground momentum state occupancy is probably a finite size effect.
At lower temperatures it can be seen that the occupancy in the quantum liquid is much larger than for the classical liquid. This is of course due to the role of the occupation entropy on the transition probability. However, the occupancy of the ground state in the quantum liquid is a small fraction of the total number of bosons in the subsystem. These fractions decrease with increasing subsystem size. Obviously this means that ground state condensation cannot account for the $\lambda$-transition or for superfluidity.
@@
Figure 2 gives the fraction of condensed bosons in the subsystem. Condensed bosons were defined as those in multiply occupied states, with a threshold set at 2 for $f_{0}$ and 3 for $f_{00}$. In the quantum liquid at the lowest temperature studied about 74% of the bosons are in states with two or more, and about 61% are in states with three or more. In contrast the fraction for the classical liquid is 51% and 23%, respectively. These specific results are for $N=1,000$, but other simulations show that these fractions are quite insensitive to the system size. The conclusion is that at these temperatures Bose-Einstein condensation is substantial, and that multiple momentum states are multiply occupied. That the majority of the bosons in the system can be considered to be condensed explains the macroscopic nature of the $\lambda$-transition and superfluidity.
It can be seen that at higher temperatures the condensation in the quantum liquid is approaching that in the classical liquid,. However even at the highest temperature studied, $T^{*}=1.00$, there is still excess condensation in the quantum liquid, $f_{0}^{\mathrm{qu}}=44\%$ compared to $f_{0}^{\mathrm{cl}}=25\%$. That there remains condensation in the quantum liquid well-above the superfluid transition temperature is likely due to the neglect in the present calculations of position permutation loops, which suppress condensation (Attard 2025 §3.2).
The kinetic energy per boson in the classical liquid is $\beta{\cal K}/N=1.4513(4)$ at $T^{*}=1.00$ and 1.4048(2) at $T^{*}=0.60$. The equipartition theorem gives the exact classical value as 3/2. Clearly the present stochastic equations of motion that use the transition probability for quantized momentum are close to the continuum classical equations of motion. The discrepancy is probably an effect of finite size. The kinetic energy per boson in the quantum liquid is $\beta{\cal K}/N=1.2248(5)$ at $T^{*}=1.00$ and 0.778(2) at $T^{*}=0.60$. The decrease in kinetic energy with decreasing temperature is a manifestation of the increasing condensation in the quantum liquid that preferentially occurs in the low lying momentum states.
@@
Figure 3 shows the shear viscosity time function for the quantum and classical liquids at the relatively high temperature of $T^{*}=1.00$. The general features of the curves are that they rise from zero at $t^{*}=0$ to reach a maximum or a plateau, in this case at about $t^{*}=6$. This maximum value is taken to be ‘the’ shear viscosity. In this and lower temperatures the quantum viscosity is rather flat by $t^{*}=6$; compare $\eta^{\mathrm{qu}}(6)=11.8(26)$ with the extrapolated maximum $\eta^{\mathrm{qu}}(10.9)=13.0$. At this highest temperature the same applies to the classical viscosity; compare $\eta^{\mathrm{cl}}(6)=11.1(32)$ with the extrapolated maximum $\eta^{\mathrm{cl}}(7.3)=11.3(34)$.
It can be seen that the viscosities of the quantum and classical liquids are statistically indistinguishable at this highest temperature studied. The fraction of condensed bosons in states with two or more bosons at this temperature is 44% for the quantum liquid and 25% for the classical liquid (Fig. 2). The average occupancy of momentum states containing two or more bosons in the quantum liquid was $\overline{N}^{\mathrm{qu,0}}_{\mathrm{occ}}=2.6258(5)$, and in the classical liquid it was $\overline{N}^{\mathrm{cl,0}}_{\mathrm{occ}}=2.1640(4)$. (The average occupancy of occupied states was $\overline{N}^{\mathrm{qu}}_{\mathrm{occ}}=1.3712(3)$, and $\overline{N}^{\mathrm{cl}}_{\mathrm{occ}}=1.1557(1)$, respectively.) Evidently at this temperature the occupation entropy has little effect on the occupation of momentum states. And the reduction in the rate of change of momentum due to occupancy has little effect on the viscosity.
The diffusion in the system was relatively small. The root mean square change in separation of a randomly chosen pair of bosons over the course of a run was 0.83(23) for the quantum liquid and 0.63(7) for the classical liquid. The statistical errors in the average potential energy and pressure were larger when estimated between runs rather than within a run. (Each run was started from an independently equilibrated configuration.) This suggests that the subsystem is somewhat glassy.